OpenStudy (anonymous):

| x -1| -2 =| 2x -5 | solve please guys help

OpenStudy (jack1):

i dont think this one is solvable dude, there's no number that satisfies the equation

OpenStudy (anonymous):

ok so what should I write

OpenStudy (jack1):

\[| x -1| -2 \neq | 2x -5 | \]

OpenStudy (jack1):

or "false" or does not equal or there is no number (positive or negative) where this works

OpenStudy (jack1):

anything where they know u tried and found the equation to be untrue will be fine dude

OpenStudy (anonymous):

ok thank you so so much :D :D

OpenStudy (jack1):

thats alright... dude to solve modulus equations it may be helpful to think of them in this form...?

OpenStudy (jack1):

| x -1| = sqrt { (x-1)^2 }

OpenStudy (jack1):

\[| x -1| = \sqrt (x-1)^2 \]

OpenStudy (anonymous):

ok thank you for everything :D